OpenStudy (anonymous):

Okay I just need someone to check my work because the site I do school on is telling me that my answer is wrong. Here is the question: Find the lateral area, total area, and volume for the cylinder with the given measurement. r = 1, h = 1 Hereis my work: LA= 2pirh LA=2pi(1)(1) LA=2pi(2) LA=4pi TA=2pirh+2pir*2 TA=4pi+2pi(1)*2 TA=4pi+2pi(1) TA=4pi+2pi TA=6pi V=pir*2h V=pi(1)*2(1) V=pi(1)(1) V-2pi What am I doing worng!?! :(

OpenStudy (anonymous):

The volume is \[\pi\cdot r^2 \cdot h\]

OpenStudy (anonymous):

I have the formula and the answer I got was 2piis this wrong???

OpenStudy (anonymous):

If r is 1 and h is 1, than\[\pi \cdot r^2 \cdot h = \pi \cdot 1^2 \cdot 1 = \pi \cdot 1 \cdot 1 = \pi\]

OpenStudy (dumbcow):

you did r*2 not r squared

OpenStudy (dumbcow):

for LA you made 1*1 = 2 ?

OpenStudy (anonymous):

V=3.14(r)squared(h) V=3.14(1)squared(1) V=3.14(1)(1) V=2pi

OpenStudy (anonymous):

I used the* sign as my square sign! I am sorry I did not tell U guys!!

OpenStudy (anonymous):

\[1\cdot1\neq2\]\[1\cdot1=1\]

OpenStudy (anonymous):

\[\pi \cdot 1 \cdot 1 \neq2 \pi\]\[\pi \cdot 1 \cdot 1 = \pi\]

OpenStudy (anonymous):

THANK YOU!! I understand what I was doing wrong now!!